Department of Mathematics, Josai University Fourth Announcement September 12th, 1996

This is the fourth announcement of the symposium

Surgery and Geometric Topology

Sept. 17 – 21, 1996

organized by

A. Ranicki and M. Yamasaki

at

Josai University, Sakado, Japan

This symposium is supported by Josai University and also and Grant-in-Aid for Scientific Research (A) of The Ministry of Education, Science, Sports and Culture of Japan.

Schedule of Talks:

September 17 (Tue)

10:00 – 11:20 Andrew Ranicki (Univ. of Edinburgh)

Chain Duality.

13:20 – 14:40 Erik Pedersen (SUNY at Binghamton)

Simplicial and Continuous Control

15:00 – 16:20 Toshiyuki Akita (Fukuoka Univ.) Cohomology and Euler Characteristics of Coxeter Groups

September 18 (Wed)

10:00 – 11:20 Frank Quinn (Virginia Tech.)

Problems with Surgery and Handlebodies in Low Dimensions

13:20 – 14:40 John Roe (Oxford Univ.)

Surgery and Operator Algebras

15:00 – 16:20 Tsuyoshi Kato (Kyoto Univ.)

Asymptotic Method on the Novikov Conjecture

16:30 – 17:20 Eiji Ogasa (Tokyo Univ.)

On the Intersection of Spheres in a Sphere

17:30 – 19:30 Reception

September 19 (Thu)

10:00 – 11:20 Bruce Hughes (Vanderbilt Univ.)

Stratified Spaces and Approximate Fibrations

13:20 – 14:40 Francis Connolly (Univ. of Notre Dame) An End Theorem for Manifold Stratified Spaces 15:00 – 16:20 Tatsuhiko Yagasaki (Kyoto Inst. of Tech.)

Infinite-dimensional Manifold Triples of Homeomorphism Groups

16:30 – 17:20 Steven Hutt (Seoul National University) A Sheaf-Theoretic Proof of the Topological Invariance of the Rational Pontrjagin Classes

September 20 (Fri)

10:00 – 11:20 Yongjin Song (Inha Univ.)

The Braid Structure of Mapping Class Groups

13:20 – 14:40 Masaharu Morimoto (Okayama Univ.)

On Fixed Point Data of Smooth Actions on Spheres

15:00 – 16:20 Masayuki Yamasaki (Josai Univ.) Epsilon Control and Perl

September 21 (Sat)

8:00 – 15:00 excursion

All lectures will be given on the 9th floor of Mizuta Memorial Library. Any changes will be announced in the following conference home page.

http://math.josai.ac.jp/~yamasaki/conference.html

Enquiries should be sent to:

Masayuki Yamasaki Department of Mathematics, Josai University 1-1 Keyaki-dai, Sakado, Saitama 350-02, Japan Tel:81-492-71-7953, Fax:81-492-71-7985

yamasaki@math.josai.ac.jp
d34494@m-unix.cc.u-tokyo.ac.jp
http://www.josai.ac.jp/home/yamasaki/